### v1

# Unbiased Estimation of Central Moments in Unbalanced Two- and Three-Level Models

Abstract
This paper derives closed-form unbiased estimators of central moments in multilevel random-effects models with unbalanced group sizes. In a two-level model, we provide unbiased estimators for the second, third, and fourth central moments under both group-level and observation-level averaging. In a three-level model, we provide unbiased estimators for the second and third central moments.
Keywords: Multilevel (hierarchical) models; unbalanced panels; variances; third central moments; fourth central moments.
JEL Classification: C13, C23
1 Introduction
Multilevel (hierarchical) linear models are widely used in empirical economics, appearing in student–teacher–school assessments and neighborhood–city–region housing markets. While estimation of variance components in multilevel random-effects models is well established, including for unbalanced designs (e.g., Searle et al., 1992; Patterson and Thompson, 1971), there appear to be few closed-form results for finite-sample unbiased estimation of third and fourth central moments. Early ANOVA and cumulant work derives unbiased estimators for certain higher-order component cumulants in variance-component settings, often in balanced designs (Tukey, 1957). More recent work studies third and fourth moment estimation focusing on consistency and asymptotic efficiency rather than explicit finite-sample unbiased within–between decompositions under alternative averaging schemes (Wu and Zhu, 2010; Wu et al., 2012).
In many economic applications, the third and fourth central moments of latent components (and their standardized counterparts, skewness and kurtosis) contain structural information about efficiency, selection, or risk. For example, work on earnings dynamics finds that income shocks exhibit substantial left-skewness and excess kurtosis (Guvenen et al., 2014). More broadly, in consumption/income risk problems and insurance markets, higher moments of persistent and transitory shocks affect welfare, precautionary saving, and pricing of downside risk (Kimball, 1990; Menezes et al., 1980). In firm and establishment data, the distribution of growth and productivity shocks is often non-Gaussian, and higher moments are informative about investment, default risk, and aggregation (Bottazzi and Secchi, 2003; Gabaix, 2011). The estimators in this paper are used in Ben-Moshe and Genesove (2026a) in a two-level model and Ben-Moshe and Genesove (2026b) in a three-level model.
This paper derives closed-form unbiased estimators of central moments in unbalanced two- and three-level models under two averaging schemes: group-level averaging (weighting clusters equally) and observation-level averaging (weighting units equally). In unbalanced panels, failing to account for unequal subgroup sizes can bias moment estimators for the latent components. Group-level averaging is natural when the sampling unit is the group, whereas observation-level averaging is natural when the sampling unit is the individual or when larger groups should receive proportionally greater weight.
Section 2 presents the two-level results (second–fourth moments), Section 3 presents the three-level results (second–third moments), and the Appendix provides the algebraic derivations.
2 Two-level model
Suppose we observe $y_{ij}$ in the two-level model
$y_{ij}=u_{i}+v_{ij},\quad i=1,\dots,n,\quad j=1,\dots,J_{i},$
where $\{u_{i}\}_{i=1}^{n}$ and $\{v_{ij}\}_{i,j}$ are unobserved, mutually independent, and within each level i.i.d., with $\mathbb{E}[u]=\mathbb{E}[v]=0$. We assume $n\geq 3$ and $J_{i}\geq 3$ for all $i$. Let $\mu_{ku}=\mathbb{E}[u^{k}]$ and $\mu_{kv}=\mathbb{E}[v^{k}]$ for $k=2,3,4$, and define sample means (superscripts $\mathrm{g}$ and $\mathrm{o}$ denote group-level and observation-level averaging) as follows,
$\bar{y}_{i}=\frac{1}{J_{i}}\sum_{j}y_{ij},\qquad\bar{y}^{\mathrm{g}}=\frac{1}{n}\sum_{i}\bar{y}_{i},\qquad\bar{y}^{\mathrm{o}}=\frac{1}{N}\sum_{i,j}y_{ij},$
where $N=\sum_{i}J_{i}$ (we suppress summation limits where the range is clear). So $J_{i}$ is the size of group $i$, $n$ is the number of groups, and $N$ is the total number of observations. Then, $\bar{y}_{i}=u_{i}+\bar{v}_{i}$, $\bar{y}^{\mathrm{g}}=\bar{u}^{\mathrm{g}}+\bar{v}^{\mathrm{g}}$, and $\bar{y}^{\mathrm{o}}=\bar{u}^{\mathrm{o}}+\bar{v}^{\mathrm{o}}$.
Unbiased estimators of $\mu_{2v}$, $\mu_{3v}$, $\mu_{4v}$ are
$\widehat{\mu}_{2v}$ $=\frac{\sum_{i,j}(y_{ij}-\bar{y}_{i})^{2}}{\sum_{i}(J_{i}-1)},\quad\widehat{\mu}_{3v}=\frac{\sum_{i,j}(y_{ij}-\bar{y}_{i})^{3}}{\sum_{i}\frac{(J_{i}-1)(J_{i}-2)}{J_{i}}},\quad\widehat{\mu}_{4v}=\frac{a_{22}^{v}\sum_{i,j}(y_{ij}-\bar{y}_{i})^{4}-a_{12}^{v}\sum_{i}\sum_{1\leq j<j^{\prime}\leq J_{i}}(y_{ij}-\bar{y}_{i})^{2}(y_{ij^{\prime}}-\bar{y}_{i})^{2}}{a_{11}^{v}a_{22}^{v}-a_{12}^{v}a_{21}^{v}}.$
The coefficients $a_{11}^{v},a_{12}^{v},a_{21}^{v},a_{22}^{v}$ are given in Appendix B.
Unbiased estimators of $\mu_{2u}$, $\mu_{3u}$, and $\mu_{4u}$ under group-level averaging are
$\widehat{\mu}_{2u}^{\mathrm{g}}$ $=\frac{1}{n-1}\sum_{i}(\bar{y}_{i}-\bar{y}^{\mathrm{g}})^{2}-\frac{\widehat{\mu}_{2v}}{n}\sum_{i}\frac{1}{J_{i}},\qquad\widehat{\mu}_{3u}^{\mathrm{g}}=\frac{n}{(n-1)(n-2)}\sum_{i}(\bar{y}_{i}-\bar{y}^{\mathrm{g}})^{3}-\frac{\widehat{\mu}_{3v}}{n}\sum_{i}\frac{1}{J_{i}^{2}},$
$\widehat{\mu}_{4u}^{\mathrm{g}}$ $=\frac{a_{22}^{\mathrm{g}}\Big(\sum_{i}(\bar{y}_{i}-\bar{y}^{\mathrm{g}})^{4}-\widehat{T}_{4}^{\mathrm{g}}\Big)-a_{12}^{\mathrm{g}}\Big(\sum_{i\neq i^{\prime}}(\bar{y}_{i}-\bar{y}^{\mathrm{g}})^{2}(\bar{y}_{i^{\prime}}-\bar{y}^{\mathrm{g}})^{2}-\widehat{T}_{22}^{\mathrm{g}}\Big)}{a_{11}^{\mathrm{g}}a_{22}^{\mathrm{g}}-a_{12}^{\mathrm{g}}a_{21}^{\mathrm{g}}}.$
The coefficients $a_{11}^{\mathrm{g}},a_{12}^{\mathrm{g}},a_{21}^{\mathrm{g}},$ and $a_{22}^{\mathrm{g}}$ and the adjustment terms $\widehat{T}_{4}^{\mathrm{g}}$ and $\widehat{T}_{22}^{\mathrm{g}}$ are given in Appendix B.
Unbiased estimators of $\mu_{2u}$, $\mu_{3u}$, and $\mu_{4u}$ under observation-level averaging are
$\widehat{\mu}_{2u}^{\mathrm{o}}$ $=\frac{\sum_{i}(\bar{y}_{i}-\bar{y}^{\mathrm{o}})^{2}-\widehat{\mu}_{2v}\sum_{i}\Big(\frac{1}{J_{i}}-\frac{1}{N}\Big)}{\sum_{i}\Big(1-\frac{2J_{i}}{N}+\frac{1}{N^{2}}\sum_{m=1}^{n}J_{m}^{2}\Big)},\quad\widehat{\mu}_{3u}^{\mathrm{o}}=\frac{\sum_{i}(\bar{y}_{i}-\bar{y}^{\mathrm{o}})^{3}-\widehat{\mu}_{3v}\sum_{i}\frac{(N-J_{i})(N-2J_{i})}{N^{2}J_{i}^{2}}}{\sum_{i}\Big(1-\frac{3J_{i}}{N}+\frac{3J_{i}^{2}}{N^{2}}-\frac{1}{N^{3}}\sum_{m=1}^{n}J_{m}^{3}\Big)},$
$\widehat{\mu}_{4u}^{\mathrm{o}}$ $=\frac{a_{22}^{\mathrm{o}}\Big(\sum_{i}(\bar{y}_{i}-\bar{y}^{\mathrm{o}})^{4}-\widehat{T}_{4}^{\mathrm{o}}\Big)-a_{12}^{\mathrm{o}}\Big(\sum_{i\neq i^{\prime}}(\bar{y}_{i}-\bar{y}^{\mathrm{o}})^{2}(\bar{y}_{i^{\prime}}-\bar{y}^{\mathrm{o}})^{2}-\widehat{T}_{22}^{\mathrm{o}}\Big)}{a_{11}^{\mathrm{o}}a_{22}^{\mathrm{o}}-a_{12}^{\mathrm{o}}a_{21}^{\mathrm{o}}}.$
The coefficients $a_{11}^{\mathrm{o}},a_{12}^{\mathrm{o}},a_{21}^{\mathrm{o}}$, and $a_{22}^{\mathrm{o}}$ and the adjustment terms $\widehat{T}_{4}^{\mathrm{o}}$ and $\widehat{T}_{22}^{\mathrm{o}}$ are given in Appendix B.
3 Three-level model
Suppose we observe $y_{ijk}$ in the three-level model,
$y_{ijk}=u_{i}+v_{ij}+w_{ijk},\quad i=1,\dots,n,\quad j=1,\dots,J_{i},\quad k=1,\dots,K_{ij},$
where $\{u_{i}\}_{i=1}^{n}$, $\{v_{ij}\}_{i,j}$, and $\{w_{ijk}\}_{i,j,k}$ are unobserved, mutually independent, and within each level i.i.d., with $\mathbb{E}[u]=\mathbb{E}[v]=\mathbb{E}[w]=0$. We assume $n\geq 3$, $J_{i}\geq 3$, and $K_{ij}\geq 3$ for all $(i,j)$. Let $\mu_{ku}=\mathbb{E}[u^{k}]$, $\mu_{kv}=\mathbb{E}[v^{k}]$, and $\mu_{kw}=\mathbb{E}[w^{k}]$ for $k=2,3$, and define sample means as follows,
$\bar{y}_{ij}$ $=\frac{1}{K_{ij}}\sum_{k}y_{ijk},\quad\bar{y}_{i}^{\mathrm{g}}=\frac{1}{J_{i}}\sum_{j}\bar{y}_{ij},\quad\bar{y}_{i}^{\mathrm{o}}=\frac{1}{K_{i}}\sum_{j}K_{ij}\bar{y}_{ij},\quad\bar{y}^{\mathrm{g}}=\frac{1}{n}\sum_{i}\bar{y}_{i}^{\mathrm{g}},\quad\bar{y}^{\mathrm{o}}=\frac{1}{N}\sum_{i,j,k}y_{ijk},$
$\bar{u}^{\mathrm{g}}$ $=\frac{1}{n}\sum_{i}u_{i},\quad\bar{u}^{\mathrm{o}}=\frac{1}{N}\sum_{i}K_{i}u_{i},\quad\bar{v}_{i}^{\mathrm{g}}=\frac{1}{J_{i}}\sum_{j}v_{ij},\quad\bar{v}_{i}^{\mathrm{o}}=\frac{1}{K_{i}}\sum_{j}K_{ij}v_{ij},\quad\bar{v}^{\mathrm{g}}=\frac{1}{n}\sum_{i}\bar{v}_{i}^{\mathrm{g}},\quad\bar{v}^{\mathrm{o}}=\frac{1}{N}\sum_{i,j}K_{ij}v_{ij},$
$\bar{w}_{ij}$ $=\frac{1}{K_{ij}}\sum_{k}w_{ijk},\quad\bar{w}_{i}^{\mathrm{g}}=\frac{1}{J_{i}}\sum_{j}\bar{w}_{ij},\quad\bar{w}_{i}^{\mathrm{o}}=\frac{1}{K_{i}}\sum_{j}K_{ij}\bar{w}_{ij},\quad\bar{w}^{\mathrm{g}}=\frac{1}{n}\sum_{i}\bar{w}_{i}^{\mathrm{g}},\quad\bar{w}^{\mathrm{o}}=\frac{1}{N}\sum_{i,j,k}w_{ijk}.$
where $K_{i}=\sum_{j}K_{ij}$ and $N=\sum_{i}K_{i}$. So $K_{ij}$ is the size of group $(i,j)$, $J_{i}$ is the size of group $i$, $K_{i}$ is the number of groups in $i$, $n$ is the number of groups overall, and $N$ is the total number of observations. Then $\bar{y}_{ij}=u_{i}+v_{ij}+\bar{w}_{ij},$  $\bar{y}_{i}^{\mathrm{g}}=u_{i}+\bar{v}_{i}^{\mathrm{g}}+\bar{w}_{i}^{\mathrm{g}},$  $\bar{y}_{i}^{\mathrm{o}}=u_{i}+\bar{v}_{i}^{\mathrm{o}}+\bar{w}_{i}^{\mathrm{o}},$  $\bar{y}^{\mathrm{g}}=\bar{u}^{\mathrm{g}}+\bar{v}^{\mathrm{g}}+\bar{w}^{\mathrm{g}},$ and $\bar{y}^{\mathrm{o}}=\bar{u}^{\mathrm{o}}+\bar{v}^{\mathrm{o}}+\bar{w}^{\mathrm{o}}.$
Unbiased estimators of $\mu_{2w}$ and $\mu_{3w}$ are
$\widehat{\mu}_{2w}$ $=\frac{\sum_{i,j,k}(y_{ijk}-\bar{y}_{ij})^{2}}{\sum_{i,j}(K_{ij}-1)},\qquad\widehat{\mu}_{3w}=\frac{\sum_{i,j,k}(y_{ijk}-\bar{y}_{ij})^{3}}{\sum_{i,j}\frac{(K_{ij}-1)(K_{ij}-2)}{K_{ij}}}.$
Unbiased estimators of $\mu_{2v}$, $\mu_{3v}$, $\mu_{2u}$, and $\mu_{3u}$ under group-level averaging are
$\widehat{\mu}_{2v}^{\mathrm{g}}$ $=\frac{\sum_{i,j}\big(\bar{y}_{ij}-\bar{y}_{i}^{\mathrm{g}}\big)^{2}-\widehat{\mu}_{2w}\sum_{i}\frac{J_{i}-1}{J_{i}}\sum_{j}\frac{1}{K_{ij}}}{\sum_{i}(J_{i}-1)},\qquad\widehat{\mu}_{3v}^{\mathrm{g}}=\frac{\sum_{i,j}\big(\bar{y}_{ij}-\bar{y}_{i}^{\mathrm{g}}\big)^{3}-\widehat{\mu}_{3w}\sum_{i}\frac{(J_{i}-1)(J_{i}-2)}{J_{i}^{2}}\sum_{j}\frac{1}{K_{ij}^{2}}}{\sum_{i}\frac{(J_{i}-1)(J_{i}-2)}{J_{i}}},$
$\widehat{\mu}_{2u}^{\mathrm{g}}$ $=\frac{1}{n-1}\sum_{i}\big(\bar{y}_{i}^{\mathrm{g}}-\bar{y}^{\mathrm{g}}\big)^{2}-\frac{\widehat{\mu}_{2v}^{\mathrm{g}}}{n}\sum_{i}\frac{1}{J_{i}}-\frac{\widehat{\mu}_{2w}}{n}\sum_{i}\frac{1}{J_{i}^{2}}\sum_{j}\frac{1}{K_{ij}},$
$\widehat{\mu}_{3u}^{\mathrm{g}}$ $=\frac{n}{(n-1)(n-2)}\sum_{i}\big(\bar{y}_{i}^{\mathrm{g}}-\bar{y}^{\mathrm{g}}\big)^{3}-\frac{\widehat{\mu}_{3v}^{\mathrm{g}}}{n}\sum_{i}\frac{1}{J_{i}^{2}}-\frac{\widehat{\mu}_{3w}}{n}\sum_{i}\frac{1}{J_{i}^{3}}\sum_{j}\frac{1}{K_{ij}^{2}}.$
Unbiased estimators of $\mu_{2v}$, $\mu_{3v}$, $\mu_{2u}$, and $\mu_{3u}$ under observation-level averaging are
$\widehat{\mu}_{2v}^{\mathrm{o}}$ $=\frac{\sum_{i,j}\big(\bar{y}_{ij}-\bar{y}_{i}^{\mathrm{o}}\big)^{2}-\widehat{\mu}_{2w}\sum_{i,j}\Big(\frac{1}{K_{ij}}-\frac{1}{K_{i}}\Big)}{\sum_{i,j}\Big(1-\frac{2K_{ij}}{K_{i}}+\frac{1}{K_{i}^{2}}\sum_{j^{\prime}=1}^{J_{i}}K_{ij^{\prime}}^{2}\Big)},\qquad\widehat{\mu}_{3v}^{\mathrm{o}}=\frac{\sum_{i,j}\big(\bar{y}_{ij}-\bar{y}_{i}^{\mathrm{o}}\big)^{3}-\widehat{\mu}_{3w}\sum_{i,j}\frac{(K_{i}-K_{ij})(K_{i}-2K_{ij})}{K_{i}^{2}K_{ij}^{2}}}{\sum_{i,j}\Big(1-\frac{3K_{ij}}{K_{i}}+\frac{3K_{ij}^{2}}{K_{i}^{2}}-\frac{1}{K_{i}^{3}}\sum_{j^{\prime}=1}^{J_{i}}K_{ij^{\prime}}^{3}\Big)},$
$\widehat{\mu}_{2u}^{\mathrm{o}}$ $=\frac{\sum_{i}\big(\bar{y}_{i}^{\mathrm{o}}-\bar{y}^{\mathrm{o}}\big)^{2}-\widehat{\mu}_{2v}^{\mathrm{o}}\sum_{i}\Bigg[\Big(1-\frac{K_{i}}{N}\Big)^{2}\frac{1}{K_{i}^{2}}\sum_{j}K_{ij}^{2}+\frac{1}{N^{2}}\sum_{i^{\prime}\neq i}\sum_{j=1}^{J_{i^{\prime}}}K_{i^{\prime}j}^{2}\Bigg]-\widehat{\mu}_{2w}\sum_{i}\Big(\frac{1}{K_{i}}-\frac{1}{N}\Big)}{\sum_{i}\Big(1-\frac{2K_{i}}{N}+\frac{1}{N^{2}}\sum_{i^{\prime}=1}^{n}K_{i^{\prime}}^{2}\Big)},$
$\widehat{\mu}_{3u}^{\mathrm{o}}$ $=\frac{\sum_{i}\big(\bar{y}_{i}^{\mathrm{o}}-\bar{y}^{\mathrm{o}}\big)^{3}-\widehat{\mu}_{3v}^{\mathrm{o}}\sum_{i}\Bigg[\Big(1-\frac{K_{i}}{N}\Big)^{3}\frac{1}{K_{i}^{3}}\sum_{j}K_{ij}^{3}+\Big(-\frac{1}{N}\Big)^{3}\sum_{i^{\prime}\neq i}\sum_{j=1}^{J_{i^{\prime}}}K_{i^{\prime}j}^{3}\Bigg]-\widehat{\mu}_{3w}\sum_{i}\frac{(N-K_{i})(N-2K_{i})}{N^{2}K_{i}^{2}}}{\sum_{i}\Big(1-\frac{3K_{i}}{N}+\frac{3K_{i}^{2}}{N^{2}}-\frac{1}{N^{3}}\sum_{i^{\prime}=1}^{n}K_{i^{\prime}}^{3}\Big)}.$
4 Conclusion
This paper derives closed-form unbiased estimators for the second, third, and fourth central moments in unbalanced two-level models, and for the second and third central moments in unbalanced three-level models. We provide these results under both group-level and observation-level averaging, allowing researchers to align estimation with their sampling design and population of interest.
The approach extends to richer multilevel designs via a three-step strategy: (i) write centered sample averages as weighted sums of latent components; (ii) apply Lemma A.1 to express expectations of powers and cross-products of these averages in terms of latent moments; and (iii) solve the resulting linear system using sample analogs and previously derived lower-order estimators. Already at the fourth moment this requires systems involving cross-products, so closed-form expressions beyond the cases treated here are best derived case-by-case for the hierarchy and moments of interest.
Appendix A Preliminaries
Lemma A.1 (Expected powers of weighted sums).
Let $x_{1},\dots,x_{m}$ be i.i.d. with $\mathbb{E}[x]=0$ and $\mu_{r}=\mathbb{E}[x^{r}]$ for $r=2,3,4$. Let $S_{m}=\sum_{\ell=1}^{m}w_{\ell}x_{\ell}$. Then
$\mathbb{E}[S_{m}^{2}]$ $=\mu_{2}\sum_{\ell=1}^{m}w_{\ell}^{2},\qquad\mathbb{E}[S_{m}^{3}]=\mu_{3}\sum_{\ell=1}^{m}w_{\ell}^{3},\qquad\mathbb{E}[S_{m}^{4}]=\mu_{4}\sum_{\ell=1}^{m}w_{\ell}^{4}+3\mu_{2}^{2}\Big(\Big(\sum_{\ell=1}^{m}w_{\ell}^{2}\Big)^{2}-\sum_{\ell=1}^{m}w_{\ell}^{4}\Big).$
Proof.
By expanding and using independence with $\mathbb{E}[x]=0$, any product term in which some $x_{\ell}$ appears to the first power has expectation zero. Therefore
$\mathbb{E}[S_{m}^{2}]$ $=\sum_{\ell=1}^{m}w_{\ell}^{2}\mathbb{E}[x_{\ell}^{2}]=\mu_{2}\sum_{\ell=1}^{m}w_{\ell}^{2},\qquad\mathbb{E}[S_{m}^{3}]=\sum_{\ell=1}^{m}w_{\ell}^{3}\mathbb{E}[x_{\ell}^{3}]=\mu_{3}\sum_{\ell=1}^{m}w_{\ell}^{3},$
$\mathbb{E}[S_{m}^{4}]$ $=\sum_{\ell=1}^{m}w_{\ell}^{4}\mathbb{E}[x_{\ell}^{4}]+6\sum_{\ell<r}w_{\ell}^{2}w_{r}^{2}\mathbb{E}[x_{\ell}^{2}]\mathbb{E}[x_{r}^{2}]=\mu_{4}\sum_{\ell=1}^{m}w_{\ell}^{4}+3\mu_{2}^{2}\Big(\Big(\sum_{\ell=1}^{m}w_{\ell}^{2}\Big)^{2}-\sum_{\ell=1}^{m}w_{\ell}^{4}\Big).\qed$
In what follows, unbiased estimators are obtained by replacing expectations with sample analogs and substituting unbiased estimators for unknown moments as they appear.
Appendix B Proofs for Two-Level Model
Using within-group variation,
$y_{ij}-\bar{y}_{i}$ $=v_{ij}-\bar{v}_{i}=\Big(1-\frac{1}{J_{i}}\Big)v_{ij}-\frac{1}{J_{i}}\sum_{j^{\prime}\neq j}v_{ij^{\prime}}.$
Applying Lemma A.1 gives
$\mathbb{E}[(y_{ij}-\bar{y}_{i})^{2}]$ $=\mu_{2v}\Big[\Big(1-\frac{1}{J_{i}}\Big)^{2}+(J_{i}-1)\Big(\frac{1}{J_{i}}\Big)^{2}\Big]=\mu_{2v}\frac{J_{i}-1}{J_{i}},$
$\mathbb{E}[(y_{ij}-\bar{y}_{i})^{3}]$ $=\mu_{3v}\Big[\Big(1-\frac{1}{J_{i}}\Big)^{3}+(J_{i}-1)\Big(-\frac{1}{J_{i}}\Big)^{3}\Big]=\mu_{3v}\frac{(J_{i}-1)(J_{i}-2)}{J_{i}^{2}},$
$\mathbb{E}[(y_{ij}-\bar{y}_{i})^{4}]$ $=\mu_{4v}\Big[\Big(1-\frac{1}{J_{i}}\Big)^{4}+(J_{i}-1)\Big(\frac{1}{J_{i}}\Big)^{4}\Big]$
$\quad+3\mu_{2v}^{2}\Bigg(\Big[\Big(1-\frac{1}{J_{i}}\Big)^{2}+(J_{i}-1)\Big(\frac{1}{J_{i}}\Big)^{2}\Big]^{2}-\Big[\Big(1-\frac{1}{J_{i}}\Big)^{4}+(J_{i}-1)\Big(\frac{1}{J_{i}}\Big)^{4}\Big]\Bigg)$
$=\mu_{4v}\frac{(J_{i}-1)(J_{i}^{2}-3J_{i}+3)}{J_{i}^{3}}+3\mu_{2v}^{2}\frac{(J_{i}-1)(2J_{i}-3)}{J_{i}^{3}},$
$\mathbb{E}[(y_{ij}-\bar{y}_{i})^{2}(y_{ij^{\prime}}-\bar{y}_{i})^{2}]$ $=\mu_{4v}\Bigg[\Big(1-\frac{1}{J_{i}}\Big)^{2}\Big(\frac{1}{J_{i}}\Big)^{2}+\Big(\frac{1}{J_{i}}\Big)^{2}\Big(1-\frac{1}{J_{i}}\Big)^{2}+(J_{i}-2)\Big(\frac{1}{J_{i}}\Big)^{4}\Bigg]$
$+\mu_{2v}^{2}\Bigg\{\Bigg[\Big(1-\frac{1}{J_{i}}\Big)^{2}+(J_{i}-1)\Big(\frac{1}{J_{i}}\Big)^{2}\Bigg]^{2}+2\Bigg[(J_{i}-2)\Big(\frac{1}{J_{i}}\Big)^{2}-\Big(1-\frac{1}{J_{i}}\Big)\Big(\frac{1}{J_{i}}\Big)-\Big(\frac{1}{J_{i}}\Big)\Big(1-\frac{1}{J_{i}}\Big)\Bigg]^{2}$
$\qquad-3\Bigg[\Big(1-\frac{1}{J_{i}}\Big)^{2}\Big(\frac{1}{J_{i}}\Big)^{2}+\Big(\frac{1}{J_{i}}\Big)^{2}\Big(1-\frac{1}{J_{i}}\Big)^{2}+(J_{i}-2)\Big(\frac{1}{J_{i}}\Big)^{4}\Bigg]\Bigg\}$
$=\mu_{4v}\frac{2J_{i}-3}{J_{i}^{3}}+\mu_{2v}^{2}\frac{J_{i}^{3}-2J_{i}^{2}-3J_{i}+9}{J_{i}^{3}}.$
Summing the fourth-moment identity over $(i,j)$ and the cross-product identity over $(i,j<j^{\prime})$ gives a $2\times 2$ system. Inverting yields
$\binom{\mu_{4v}}{\mu_{2v}^{2}}=\begin{pmatrix}a_{11}^{v}&a_{12}^{v}\\ a_{21}^{v}&a_{22}^{v}\end{pmatrix}^{-1}\binom{\mathbb{E}\!\left[\sum_{i=1}^{n}\sum_{j=1}^{J_{i}}(y_{ij}-\bar{y}_{i})^{4}\right]}{\mathbb{E}\!\left[\sum_{i=1}^{n}\sum_{1\leq j<j^{\prime}\leq J_{i}}(y_{ij}-\bar{y}_{i})^{2}(y_{ij^{\prime}}-\bar{y}_{i})^{2}\right]}$,
where
$a_{11}^{v}$ $=\sum_{i=1}^{n}\frac{(J_{i}-1)(J_{i}^{2}-3J_{i}+3)}{J_{i}^{2}},\qquad a_{12}^{v}=3\sum_{i=1}^{n}\frac{(J_{i}-1)(2J_{i}-3)}{J_{i}^{2}},$
$a_{21}^{v}$ $=\frac{1}{2}\sum_{i=1}^{n}\frac{(J_{i}-1)(2J_{i}-3)}{J_{i}^{2}},\qquad a_{22}^{v}=\frac{1}{2}\sum_{i=1}^{n}\frac{(J_{i}-1)(J_{i}^{3}-2J_{i}^{2}-3J_{i}+9)}{J_{i}^{2}}.$
Using between-group variation and group-averages,
$\bar{y}_{i}-\bar{y}^{\mathrm{g}}$ $=\Big[u_{i}-\frac{1}{n}\sum_{i^{\prime}=1}^{n}u_{i^{\prime}}\Big]+\Big[\bar{v}_{i}-\frac{1}{n}\sum_{i^{\prime}=1}^{n}\bar{v}_{i^{\prime}}\Big]=\Big[\Big(1-\frac{1}{n}\Big)u_{i}-\frac{1}{n}\sum_{i^{\prime}\neq i}u_{i^{\prime}}\Big]+\Big[\Big(1-\frac{1}{n}\Big)\bar{v}_{i}-\frac{1}{n}\sum_{i^{\prime}\neq i}\bar{v}_{i^{\prime}}\Big].$
Applying Lemma A.1 and using independence gives
$\mathbb{E}[(\bar{y}_{i}-\bar{y}^{\mathrm{g}})^{2}]$ $=\mu_{2u}\Big[\Big(1-\frac{1}{n}\Big)^{2}+(n-1)\Big(\frac{1}{n}\Big)^{2}\Big]+\mu_{2v}\Big[\Big(1-\frac{1}{n}\Big)^{2}\frac{1}{J_{i}}+\Big(\frac{1}{n}\Big)^{2}\sum_{i^{\prime}\neq i}\frac{1}{J_{i^{\prime}}}\Big]$
$=\mu_{2u}\frac{n-1}{n}+\mu_{2v}\Big(\frac{1}{J_{i}}-\frac{2}{nJ_{i}}+\frac{1}{n^{2}}\sum_{i^{\prime}=1}^{n}\frac{1}{J_{i^{\prime}}}\Big),$
$\mathbb{E}[(\bar{y}_{i}-\bar{y}^{\mathrm{g}})^{3}]$ $=\mu_{3u}\Big[\Big(1-\frac{1}{n}\Big)^{3}+(n-1)\Big(-\frac{1}{n}\Big)^{3}\Big]+\mu_{3v}\Big[\Big(1-\frac{1}{n}\Big)^{3}\frac{1}{J_{i}^{2}}+\Big(-\frac{1}{n}\Big)^{3}\sum_{i^{\prime}\neq i}\frac{1}{J_{i^{\prime}}^{2}}\Big]$
$=\mu_{3u}\frac{(n-1)(n-2)}{n^{2}}+\mu_{3v}\Big(\frac{1}{J_{i}^{2}}-\frac{3}{nJ_{i}^{2}}+\frac{3}{n^{2}J_{i}^{2}}-\frac{1}{n^{3}}\sum_{i^{\prime}=1}^{n}\frac{1}{J_{i^{\prime}}^{2}}\Big),$
$\mathbb{E}[(\bar{y}_{i}-\bar{y}^{\mathrm{g}})^{4}]$ $=\mathbb{E}[(u_{i}-\bar{u}^{\mathrm{g}})^{4}]+6\,\mathbb{E}[(u_{i}-\bar{u}^{\mathrm{g}})^{2}]\mathbb{E}[(\bar{v}_{i}-\bar{v}^{\mathrm{g}})^{2}]+\mathbb{E}[(\bar{v}_{i}-\bar{v}^{\mathrm{g}})^{4}]$
$=\mu_{4u}\frac{(n-1)(n^{2}-3n+3)}{n^{3}}+3\mu_{2u}^{2}\frac{(n-1)(2n-3)}{n^{3}}$
$\quad+6\mu_{2u}\mu_{2v}\frac{n-1}{n}\Big(\frac{1}{J_{i}}-\frac{2}{nJ_{i}}+\frac{1}{n^{2}}\sum_{i^{\prime}=1}^{n}\frac{1}{J_{i^{\prime}}}\Big)+\mu_{4v}\Big[\Big(1-\frac{1}{n}\Big)^{4}\frac{1}{J_{i}^{3}}+\Big(\frac{1}{n}\Big)^{4}\sum_{i^{\prime}\neq i}\frac{1}{J_{i^{\prime}}^{3}}\Big]$
$\quad+3\mu_{2v}^{2}\Bigg(\Big[\Big(1-\frac{1}{n}\Big)^{2}\frac{1}{J_{i}}+\Big(\frac{1}{n}\Big)^{2}\sum_{i^{\prime}\neq i}\frac{1}{J_{i^{\prime}}}\Big]^{2}-\Big[\Big(1-\frac{1}{n}\Big)^{4}\frac{1}{J_{i}^{3}}+\Big(\frac{1}{n}\Big)^{4}\sum_{i^{\prime}\neq i}\frac{1}{J_{i^{\prime}}^{3}}\Big]\Bigg),$
$\mathbb{E}[(\bar{y}_{i}-\bar{y}^{\mathrm{g}})^{2}(\bar{y}_{i^{\prime}}-\bar{y}^{\mathrm{g}})^{2}]$ $=\mu_{4u}\frac{2n-3}{n^{3}}+\mu_{2u}^{2}\frac{n^{3}+6n^{2}-11n+9}{n^{3}}-4\mu_{2u}\mu_{2v}\frac{n-1}{n^{2}}\Bigg(\frac{1}{n^{2}}\sum_{m=1}^{n}\frac{1}{J_{m}}-\frac{1}{nJ_{i}}-\frac{1}{nJ_{i^{\prime}}}\Bigg)$
$\quad+\mu_{2u}\mu_{2v}\frac{n-1}{n}\Bigg[\Big(\frac{1}{J_{i}}-\frac{2}{nJ_{i}}+\frac{1}{n^{2}}\sum_{m=1}^{n}\frac{1}{J_{m}}\Big)+\Big(\frac{1}{J_{i^{\prime}}}-\frac{2}{nJ_{i^{\prime}}}+\frac{1}{n^{2}}\sum_{m=1}^{n}\frac{1}{J_{m}}\Big)\Bigg]$
$\quad+\mu_{2v}^{2}\Bigg[\Big(\frac{1}{J_{i}}-\frac{2}{nJ_{i}}+\frac{1}{n^{2}}\sum_{m=1}^{n}\frac{1}{J_{m}}\Big)\Big(\frac{1}{J_{i^{\prime}}}-\frac{2}{nJ_{i^{\prime}}}+\frac{1}{n^{2}}\sum_{m=1}^{n}\frac{1}{J_{m}}\Big)+2\Big(\frac{1}{n^{2}}\sum_{m=1}^{n}\frac{1}{J_{m}}-\frac{1}{nJ_{i}}-\frac{1}{nJ_{i^{\prime}}}\Big)^{2}\Bigg]$
$\quad+\big(\mu_{4v}-3\mu_{2v}^{2}\big)\frac{1}{n^{4}}\Bigg[\sum_{m=1}^{n}\frac{1}{J_{m}^{3}}+(n^{2}-2n)\Big(\frac{1}{J_{i}^{3}}+\frac{1}{J_{i^{\prime}}^{3}}\Big)\Bigg].$
Summing the fourth-moment identity over $i$ and the cross-product identity over $i\neq i^{\prime}$ gives a $2\times 2$ system. Inverting yields
$\binom{\mu_{4u}}{\mu_{2u}^{2}}=\begin{pmatrix}a_{11}^{\mathrm{g}}&a_{12}^{\mathrm{g}}\\ a_{21}^{\mathrm{g}}&a_{22}^{\mathrm{g}}\end{pmatrix}^{-1}\binom{\mathbb{E}\!\left[\sum_{i=1}^{n}(\bar{y}_{i}-\bar{y}^{\mathrm{g}})^{4}-T_{4}^{\mathrm{g}}\right]}{\mathbb{E}\!\left[\sum_{i\neq i^{\prime}}(\bar{y}_{i}-\bar{y}^{\mathrm{g}})^{2}(\bar{y}_{i^{\prime}}-\bar{y}^{\mathrm{g}})^{2}-T_{22}^{\mathrm{g}}\right]}$,
where
$a_{11}^{\mathrm{g}}$ $=\frac{(n-1)(n^{2}-3n+3)}{n^{2}},\quad a_{12}^{\mathrm{g}}=\frac{3(n-1)(2n-3)}{n^{2}},\quad a_{21}^{\mathrm{g}}=\frac{(n-1)(2n-3)}{n^{2}},\quad a_{22}^{\mathrm{g}}=\frac{(n-1)(n^{3}+6n^{2}-11n+9)}{n^{2}},$
$T_{4}^{\mathrm{g}}$ $=6\mu_{2u}\mu_{2v}\,\frac{(n-1)^{2}}{n^{2}}\sum_{i=1}^{n}\frac{1}{J_{i}}+\mu_{4v}\,\frac{(n-1)(n^{2}-3n+3)}{n^{3}}\sum_{i=1}^{n}\frac{1}{J_{i}^{3}}$
$\quad+3\mu_{2v}^{2}\Bigg(\frac{(n-2)^{2}}{n^{2}}\sum_{i=1}^{n}\frac{1}{J_{i}^{2}}+\frac{2n-3}{n^{3}}\Big(\sum_{i=1}^{n}\frac{1}{J_{i}}\Big)^{2}-\frac{(n-1)(n^{2}-3n+3)}{n^{3}}\sum_{i=1}^{n}\frac{1}{J_{i}^{3}}\Bigg),$
$T_{22}^{\mathrm{g}}$ $=2\mu_{2u}\mu_{2v}\,\frac{(n-1)\big((n-1)^{2}+2\big)}{n^{2}}\sum_{i=1}^{n}\frac{1}{J_{i}}+\mu_{4v}\,\frac{(n-1)(2n-3)}{n^{3}}\sum_{i=1}^{n}\frac{1}{J_{i}^{3}}$
$\quad+\mu_{2v}^{2}\Bigg(\frac{n^{3}-2n^{2}-3n+9}{n^{3}}\Big(\sum_{i=1}^{n}\frac{1}{J_{i}}\Big)^{2}+\frac{(n-2)(6-n)}{n^{2}}\sum_{i=1}^{n}\frac{1}{J_{i}^{2}}-3\,\frac{(n-1)(2n-3)}{n^{3}}\sum_{i=1}^{n}\frac{1}{J_{i}^{3}}\Bigg).$
Similarly, using between-group variation and observation-averages,
$\bar{y}_{i}-\bar{y}^{\mathrm{o}}$ $=\Big[\Big(1-\frac{J_{i}}{N}\Big)u_{i}-\frac{1}{N}\sum_{i^{\prime}\neq i}J_{i^{\prime}}u_{i^{\prime}}\Big]+\Big[\Big(\frac{1}{J_{i}}-\frac{1}{N}\Big)\sum_{j=1}^{J_{i}}v_{ij}-\frac{1}{N}\sum_{i^{\prime}\neq i}\sum_{j=1}^{J_{i^{\prime}}}v_{i^{\prime}j}\Big].$
Applying Lemma A.1 and using independence gives
$\mathbb{E}[(\bar{y}_{i}-\bar{y}^{\mathrm{o}})^{2}]$ $=\mu_{2u}\Big[\Big(1-\frac{J_{i}}{N}\Big)^{2}+\frac{1}{N^{2}}\sum_{i^{\prime}\neq i}J_{i^{\prime}}^{2}\Big]+\mu_{2v}\Big[J_{i}\Big(\frac{1}{J_{i}}-\frac{1}{N}\Big)^{2}+(N-J_{i})\Big(\frac{1}{N}\Big)^{2}\Big]$
$=\mu_{2u}\Big(1-\frac{2J_{i}}{N}+\frac{1}{N^{2}}\sum_{m=1}^{n}J_{m}^{2}\Big)+\mu_{2v}\Big(\frac{1}{J_{i}}-\frac{1}{N}\Big),$
$\mathbb{E}[(\bar{y}_{i}-\bar{y}^{\mathrm{o}})^{3}]$ $=\mu_{3u}\Big[\Big(1-\frac{J_{i}}{N}\Big)^{3}-\frac{1}{N^{3}}\sum_{i^{\prime}\neq i}J_{i^{\prime}}^{3}\Big]+\mu_{3v}\Big[J_{i}\Big(\frac{1}{J_{i}}-\frac{1}{N}\Big)^{3}-(N-J_{i})\Big(\frac{1}{N}\Big)^{3}\Big]$
$=\mu_{3u}\Big(1-\frac{3J_{i}}{N}+\frac{3J_{i}^{2}}{N^{2}}-\frac{1}{N^{3}}\sum_{m=1}^{n}J_{m}^{3}\Big)+\mu_{3v}\frac{(N-J_{i})(N-2J_{i})}{N^{2}J_{i}^{2}},$
$\mathbb{E}[(\bar{y}_{i}-\bar{y}^{\mathrm{o}})^{4}]$ $=\mathbb{E}[(u_{i}-\bar{u}^{\mathrm{o}})^{4}]+6\,\mathbb{E}[(u_{i}-\bar{u}^{\mathrm{o}})^{2}]\mathbb{E}[(\bar{v}_{i}-\bar{v}^{\mathrm{o}})^{2}]+\mathbb{E}[(\bar{v}_{i}-\bar{v}^{\mathrm{o}})^{4}]$
$=\mu_{4u}\Big[\Big(1-\frac{J_{i}}{N}\Big)^{4}+\frac{1}{N^{4}}\sum_{i^{\prime}\neq i}J_{i^{\prime}}^{4}\Big]+3\mu_{2u}^{2}\Bigg(\Big[\Big(1-\frac{J_{i}}{N}\Big)^{2}+\frac{1}{N^{2}}\sum_{i^{\prime}\neq i}J_{i^{\prime}}^{2}\Big]^{2}-\Big[\Big(1-\frac{J_{i}}{N}\Big)^{4}+\frac{1}{N^{4}}\sum_{i^{\prime}\neq i}J_{i^{\prime}}^{4}\Big]\Bigg)$
$\quad+6\mu_{2u}\mu_{2v}\Big[\Big(1-\frac{J_{i}}{N}\Big)^{2}+\frac{1}{N^{2}}\sum_{i^{\prime}\neq i}J_{i^{\prime}}^{2}\Big]\Big(\frac{1}{J_{i}}-\frac{1}{N}\Big)+\mu_{4v}\Big[J_{i}\Big(\frac{1}{J_{i}}-\frac{1}{N}\Big)^{4}+(N-J_{i})\Big(\frac{1}{N}\Big)^{4}\Big]$
$\quad+3\mu_{2v}^{2}\Bigg(\Big(\frac{1}{J_{i}}-\frac{1}{N}\Big)^{2}-\Big[J_{i}\Big(\frac{1}{J_{i}}-\frac{1}{N}\Big)^{4}+(N-J_{i})\Big(\frac{1}{N}\Big)^{4}\Big]\Bigg),$
$\mathbb{E}\!\left[(\bar{y}_{i}-\bar{y}^{\mathrm{o}})^{2}(\bar{y}_{i^{\prime}}-\bar{y}^{\mathrm{o}})^{2}\right]$ $=\mu_{4u}\Bigg[\frac{J_{i}^{2}}{N^{2}}\Big(1-\frac{J_{i}}{N}\Big)^{2}+\frac{J_{i^{\prime}}^{2}}{N^{2}}\Big(1-\frac{J_{i^{\prime}}}{N}\Big)^{2}+\frac{1}{N^{4}}\Big(\sum_{m=1}^{n}J_{m}^{4}-J_{i}^{4}-J_{i^{\prime}}^{4}\Big)\Bigg]$
$\quad+\mu_{2u}^{2}\Bigg(\Big(1-\frac{2J_{i}}{N}+\frac{1}{N^{2}}\sum_{m=1}^{n}J_{m}^{2}\Big)\Big(1-\frac{2J_{i^{\prime}}}{N}+\frac{1}{N^{2}}\sum_{m=1}^{n}J_{m}^{2}\Big)+2\Big(\frac{1}{N^{2}}\sum_{m=1}^{n}J_{m}^{2}-\frac{J_{i}+J_{i^{\prime}}}{N}\Big)^{2}$
$\qquad\qquad-3\Bigg[\frac{J_{i}^{2}}{N^{2}}\Big(1-\frac{J_{i}}{N}\Big)^{2}+\frac{J_{i^{\prime}}^{2}}{N^{2}}\Big(1-\frac{J_{i^{\prime}}}{N}\Big)^{2}+\frac{1}{N^{4}}\Big(\sum_{m=1}^{n}J_{m}^{4}-J_{i}^{4}-J_{i^{\prime}}^{4}\Big)\Bigg]\Bigg)$
$\quad+\mu_{2u}\mu_{2v}\Bigg[\Big(1-\frac{2J_{i}}{N}+\frac{1}{N^{2}}\sum_{m=1}^{n}J_{m}^{2}\Big)\Big(\frac{1}{J_{i^{\prime}}}-\frac{1}{N}\Big)+\Big(1-\frac{2J_{i^{\prime}}}{N}+\frac{1}{N^{2}}\sum_{m=1}^{n}J_{m}^{2}\Big)\Big(\frac{1}{J_{i}}-\frac{1}{N}\Big)$
$\qquad\qquad\qquad\qquad\qquad\qquad-\frac{4}{N}\Big(\frac{1}{N^{2}}\sum_{m=1}^{n}J_{m}^{2}-\frac{J_{i}+J_{i^{\prime}}}{N}\Big)\Bigg]$
$\quad+\mu_{4v}\Bigg[\frac{1}{N^{4}}\Bigg(\frac{(N-J_{i})^{2}}{J_{i}}+\frac{(N-J_{i^{\prime}})^{2}}{J_{i^{\prime}}}+(N-J_{i}-J_{i^{\prime}})\Bigg)\Bigg]$
$\quad+\mu_{2v}^{2}\Bigg(\Big(\frac{1}{J_{i}}-\frac{1}{N}\Big)\Big(\frac{1}{J_{i^{\prime}}}-\frac{1}{N}\Big)+\frac{2}{N^{2}}-3\cdot\frac{1}{N^{4}}\Bigg(\frac{(N-J_{i})^{2}}{J_{i}}+\frac{(N-J_{i^{\prime}})^{2}}{J_{i^{\prime}}}+(N-J_{i}-J_{i^{\prime}})\Bigg)\Bigg).$
Summing the fourth-moment identity over $i$ and the cross-product identity over $i\neq i^{\prime}$ gives a $2\times 2$ system.
$\binom{\mu_{4u}}{\mu_{2u}^{2}}=\begin{pmatrix}a_{11}^{\mathrm{o}}&a_{12}^{\mathrm{o}}\\ a_{21}^{\mathrm{o}}&a_{22}^{\mathrm{o}}\end{pmatrix}^{-1}\binom{\mathbb{E}\!\left[\sum_{i=1}^{n}(\bar{y}_{i}-\bar{y}^{\mathrm{o}})^{4}-T_{4}^{\mathrm{o}}\right]}{\mathbb{E}\!\left[\sum_{i\neq i^{\prime}}(\bar{y}_{i}-\bar{y}^{\mathrm{o}})^{2}(\bar{y}_{i^{\prime}}-\bar{y}^{\mathrm{o}})^{2}-T_{22}^{\mathrm{o}}\right]}$,
where
$a_{11}^{\mathrm{o}}$ $=\sum_{i=1}^{n}\Bigg[\Big(1-\frac{J_{i}}{N}\Big)^{4}+\frac{1}{N^{4}}\sum_{m\neq i}J_{m}^{4}\Bigg],\quad a_{12}^{\mathrm{o}}=3\sum_{i=1}^{n}\Bigg(\Big[\Big(1-\frac{J_{i}}{N}\Big)^{2}+\frac{1}{N^{2}}\sum_{m\neq i}J_{m}^{2}\Big]^{2}-\Big[\Big(1-\frac{J_{i}}{N}\Big)^{4}+\frac{1}{N^{4}}\sum_{m\neq i}J_{m}^{4}\Big]\Bigg),$
$a_{21}^{\mathrm{o}}$ $=\sum_{i\neq i^{\prime}}\Bigg[\Big(\frac{J_{i}}{N}\Big)^{2}\Big(1-\frac{J_{i}}{N}\Big)^{2}+\Big(\frac{J_{i^{\prime}}}{N}\Big)^{2}\Big(1-\frac{J_{i^{\prime}}}{N}\Big)^{2}+\frac{1}{N^{4}}\sum_{m\neq i,i^{\prime}}J_{m}^{4}\Bigg],$
$a_{22}^{\mathrm{o}}$ $=\sum_{i\neq i^{\prime}}\Bigg(\Big[1-\frac{2J_{i}}{N}+\frac{1}{N^{2}}\sum_{m=1}^{n}J_{m}^{2}\Big]\Big[1-\frac{2J_{i^{\prime}}}{N}+\frac{1}{N^{2}}\sum_{m=1}^{n}J_{m}^{2}\Big]+2\Big[\frac{1}{N^{2}}\sum_{m=1}^{n}J_{m}^{2}-\frac{J_{i}+J_{i^{\prime}}}{N}\Big]^{2}$
$\qquad\qquad-3\Bigg[\Big(\frac{J_{i}}{N}\Big)^{2}\Big(1-\frac{J_{i}}{N}\Big)^{2}+\Big(\frac{J_{i^{\prime}}}{N}\Big)^{2}\Big(1-\frac{J_{i^{\prime}}}{N}\Big)^{2}+\frac{1}{N^{4}}\sum_{m\neq i,i^{\prime}}J_{m}^{4}\Bigg]\Bigg),$
$T_{4}^{\mathrm{o}}$ $=6\mu_{2u}\mu_{2v}\sum_{i=1}^{n}\Big(1-\frac{2J_{i}}{N}+\frac{1}{N^{2}}\sum_{m=1}^{n}J_{m}^{2}\Big)\Big(\frac{1}{J_{i}}-\frac{1}{N}\Big)+\mu_{4v}\sum_{i=1}^{n}\Bigg[J_{i}\Big(\frac{1}{J_{i}}-\frac{1}{N}\Big)^{4}+(N-J_{i})\Big(\frac{1}{N}\Big)^{4}\Bigg]$
$\quad+3\mu_{2v}^{2}\sum_{i=1}^{n}\Bigg[\Big(\frac{1}{J_{i}}-\frac{1}{N}\Big)^{2}-\Bigg(J_{i}\Big(\frac{1}{J_{i}}-\frac{1}{N}\Big)^{4}+(N-J_{i})\Big(\frac{1}{N}\Big)^{4}\Bigg)\Bigg],$
$T_{22}^{\mathrm{o}}$ $=\mu_{2u}\mu_{2v}\sum_{i\neq i^{\prime}}\Bigg[\Big(1-\frac{2J_{i}}{N}+\frac{1}{N^{2}}\sum_{m=1}^{n}J_{m}^{2}\Big)\Big(\frac{1}{J_{i^{\prime}}}-\frac{1}{N}\Big)+\Big(1-\frac{2J_{i^{\prime}}}{N}+\frac{1}{N^{2}}\sum_{m=1}^{n}J_{m}^{2}\Big)\Big(\frac{1}{J_{i}}-\frac{1}{N}\Big)$
$\qquad\qquad\qquad\qquad-\frac{4}{N}\Big(\frac{1}{N^{2}}\sum_{m=1}^{n}J_{m}^{2}-\frac{J_{i}+J_{i^{\prime}}}{N}\Big)\Bigg]+\mu_{4v}\sum_{i\neq i^{\prime}}\frac{1}{N^{4}}\Bigg[\frac{(N-J_{i})^{2}}{J_{i}}+\frac{(N-J_{i^{\prime}})^{2}}{J_{i^{\prime}}}+(N-J_{i}-J_{i^{\prime}})\Bigg]$
$\quad+\mu_{2v}^{2}\sum_{i\neq i^{\prime}}\Bigg[\Big(\frac{1}{J_{i}}-\frac{1}{N}\Big)\Big(\frac{1}{J_{i^{\prime}}}-\frac{1}{N}\Big)+\frac{2}{N^{2}}-\frac{3}{N^{4}}\Bigg(\frac{(N-J_{i})^{2}}{J_{i}}+\frac{(N-J_{i^{\prime}})^{2}}{J_{i^{\prime}}}+(N-J_{i}-J_{i^{\prime}})\Bigg)\Bigg].$
Appendix C Proofs for Three-Level Model
Using within-$(i,j)$ variation,
$y_{ijk}-\bar{y}_{ij}$ $=w_{ijk}-\bar{w}_{ij}=\Big(1-\frac{1}{K_{ij}}\Big)w_{ijk}-\frac{1}{K_{ij}}\sum_{k^{\prime}\neq k}w_{ijk^{\prime}}.$
Applying Lemma A.1 gives
$\mathbb{E}[(y_{ijk}-\bar{y}_{ij})^{2}]$ $=\mu_{2w}\Big[\Big(1-\frac{1}{K_{ij}}\Big)^{2}+(K_{ij}-1)\Big(\frac{1}{K_{ij}}\Big)^{2}\Big]=\mu_{2w}\frac{K_{ij}-1}{K_{ij}},$
$\mathbb{E}[(y_{ijk}-\bar{y}_{ij})^{3}]$ $=\mu_{3w}\Big[\Big(1-\frac{1}{K_{ij}}\Big)^{3}+(K_{ij}-1)\Big(-\frac{1}{K_{ij}}\Big)^{3}\Big]=\mu_{3w}\frac{(K_{ij}-1)(K_{ij}-2)}{K_{ij}^{2}}.$
Using within-$i$ variation and group-averages across $j$,
$\bar{y}_{ij}-\bar{y}_{i}^{\mathrm{g}}$ $=[v_{ij}-\bar{v}_{i}^{\mathrm{g}}]+[\bar{w}_{ij}-\bar{w}_{i}^{\mathrm{g}}]=\Big[\Big(1-\frac{1}{J_{i}}\Big)v_{ij}-\frac{1}{J_{i}}\sum_{j^{\prime}\neq j}v_{ij^{\prime}}\Big]+\Big[\Big(1-\frac{1}{J_{i}}\Big)\bar{w}_{ij}-\frac{1}{J_{i}}\sum_{j^{\prime}\neq j}\bar{w}_{ij^{\prime}}\Big],$
Applying Lemma A.1 and using independence gives
$\mathbb{E}[(\bar{y}_{ij}-\bar{y}_{i}^{\mathrm{g}})^{2}]$ $=\mu_{2v}\Big[\Big(1-\frac{1}{J_{i}}\Big)^{2}+(J_{i}-1)\Big(\frac{1}{J_{i}}\Big)^{2}\Big]+\mu_{2w}\Big[\Big(1-\frac{1}{J_{i}}\Big)^{2}\frac{1}{K_{ij}}+\Big(\frac{1}{J_{i}}\Big)^{2}\sum_{j^{\prime}\neq j}\frac{1}{K_{ij^{\prime}}}\Big]$
$=\mu_{2v}\frac{J_{i}-1}{J_{i}}+\mu_{2w}\Big(\frac{1}{K_{ij}}-\frac{2}{J_{i}K_{ij}}+\frac{1}{J_{i}^{2}}\sum_{j^{\prime}=1}^{J_{i}}\frac{1}{K_{ij^{\prime}}}\Big),$
$\mathbb{E}[(\bar{y}_{ij}-\bar{y}_{i}^{\mathrm{g}})^{3}]$ $=\mu_{3v}\Big[\Big(1-\frac{1}{J_{i}}\Big)^{3}+(J_{i}-1)\Big(-\frac{1}{J_{i}}\Big)^{3}\Big]+\mu_{3w}\Big[\Big(1-\frac{1}{J_{i}}\Big)^{3}\frac{1}{K_{ij}^{2}}+\Big(-\frac{1}{J_{i}}\Big)^{3}\sum_{j^{\prime}\neq j}\frac{1}{K_{ij^{\prime}}^{2}}\Big]$
$=\mu_{3v}\frac{(J_{i}-1)(J_{i}-2)}{J_{i}^{2}}+\mu_{3w}\Big(\frac{1}{K_{ij}^{2}}-\frac{3}{J_{i}K_{ij}^{2}}+\frac{3}{J_{i}^{2}K_{ij}^{2}}-\frac{1}{J_{i}^{3}}\sum_{j^{\prime}=1}^{J_{i}}\frac{1}{K_{ij^{\prime}}^{2}}\Big).$
Similarly, using within-$i$ variation and observation-averages across $j$,
$\bar{y}_{ij}-\bar{y}_{i}^{\mathrm{o}}$ $=[v_{ij}-\bar{v}_{i}^{\mathrm{o}}]+[\bar{w}_{ij}-\bar{w}_{i}^{\mathrm{o}}]=\Big[\Big(1-\frac{K_{ij}}{K_{i}}\Big)v_{ij}-\frac{1}{K_{i}}\sum_{j^{\prime}\neq j}K_{ij^{\prime}}v_{ij^{\prime}}\Big]+\Big[\Big(1-\frac{K_{ij}}{K_{i}}\Big)\bar{w}_{ij}-\frac{1}{K_{i}}\sum_{j^{\prime}\neq j}K_{ij^{\prime}}\bar{w}_{ij^{\prime}}\Big].$
Applying Lemma A.1 and using independence gives
$\mathbb{E}[(\bar{y}_{ij}-\bar{y}_{i}^{\mathrm{o}})^{2}]$ $=\mu_{2v}\Big[\Big(1-\frac{K_{ij}}{K_{i}}\Big)^{2}+\frac{1}{K_{i}^{2}}\sum_{j^{\prime}\neq j}K_{ij^{\prime}}^{2}\Big]+\mu_{2w}\Big[\Big(1-\frac{K_{ij}}{K_{i}}\Big)^{2}\frac{1}{K_{ij}}+\frac{1}{K_{i}^{2}}\sum_{j^{\prime}\neq j}\frac{K_{ij^{\prime}}^{2}}{K_{ij^{\prime}}}\Big]$
$=\mu_{2v}\Big(1-\frac{2K_{ij}}{K_{i}}+\frac{1}{K_{i}^{2}}\sum_{j^{\prime}=1}^{J_{i}}K_{ij^{\prime}}^{2}\Big)+\mu_{2w}\Big(\frac{1}{K_{ij}}-\frac{1}{K_{i}}\Big),$
$\mathbb{E}[(\bar{y}_{ij}-\bar{y}_{i}^{\mathrm{o}})^{3}]$ $=\mu_{3v}\Big[\Big(1-\frac{K_{ij}}{K_{i}}\Big)^{3}-\frac{1}{K_{i}^{3}}\sum_{j^{\prime}\neq j}K_{ij^{\prime}}^{3}\Big]+\mu_{3w}\Big[\Big(1-\frac{K_{ij}}{K_{i}}\Big)^{3}\frac{1}{K_{ij}^{2}}-\frac{1}{K_{i}^{3}}\sum_{j^{\prime}\neq j}\frac{K_{ij^{\prime}}^{3}}{K_{ij^{\prime}}^{2}}\Big]$
$=\mu_{3v}\Big(1-\frac{3K_{ij}}{K_{i}}+\frac{3K_{ij}^{2}}{K_{i}^{2}}-\frac{1}{K_{i}^{3}}\sum_{j^{\prime}=1}^{J_{i}}K_{ij^{\prime}}^{3}\Big)+\mu_{3w}\frac{(K_{i}-K_{ij})(K_{i}-2K_{ij})}{K_{i}^{2}K_{ij}^{2}}.$
Using between-$i$ variation and group-averages
$\bar{y}_{i}^{\mathrm{g}}-\bar{y}^{\mathrm{g}}=[u_{i}-\bar{u}^{\mathrm{g}}]+[\bar{v}_{i}^{\mathrm{g}}-\bar{v}^{\mathrm{g}}]+[\bar{w}_{i}^{\mathrm{g}}-\bar{w}^{\mathrm{g}}]$ $=\Big[\Big(1-\frac{1}{n}\Big)u_{i}-\frac{1}{n}\sum_{i^{\prime}\neq i}u_{i^{\prime}}\Big]+\Bigg[\Big(1-\frac{1}{n}\Big)\frac{1}{J_{i}}\sum_{j=1}^{J_{i}}v_{ij}-\frac{1}{n}\sum_{i^{\prime}\neq i}\frac{1}{J_{i^{\prime}}}\sum_{j=1}^{J_{i^{\prime}}}v_{i^{\prime}j}\Bigg]$
$\quad+\Bigg[\Big(1-\frac{1}{n}\Big)\frac{1}{J_{i}}\sum_{j=1}^{J_{i}}\frac{1}{K_{ij}}\sum_{k=1}^{K_{ij}}w_{ijk}-\frac{1}{n}\sum_{i^{\prime}\neq i}\frac{1}{J_{i^{\prime}}}\sum_{j=1}^{J_{i^{\prime}}}\frac{1}{K_{i^{\prime}j}}\sum_{k=1}^{K_{i^{\prime}j}}w_{i^{\prime}jk}\Bigg].$
Applying Lemma A.1 and using independence gives
$\mathbb{E}[(\bar{y}_{i}^{\mathrm{g}}-\bar{y}^{\mathrm{g}})^{2}]$ $=\mu_{2u}\frac{n-1}{n}+\mu_{2v}\Big(\frac{1}{J_{i}}-\frac{2}{nJ_{i}}+\frac{1}{n^{2}}\sum_{i^{\prime}=1}^{n}\frac{1}{J_{i^{\prime}}}\Big)+\mu_{2w}\Bigg(\Big(1-\frac{1}{n}\Big)^{2}\frac{1}{J_{i}^{2}}\sum_{j=1}^{J_{i}}\frac{1}{K_{ij}}+\Big(\frac{1}{n}\Big)^{2}\sum_{i^{\prime}\neq i}\frac{1}{J_{i^{\prime}}^{2}}\sum_{j=1}^{J_{i^{\prime}}}\frac{1}{K_{i^{\prime}j}}\Bigg),$
$\mathbb{E}[(\bar{y}_{i}^{\mathrm{g}}-\bar{y}^{\mathrm{g}})^{3}]$ $=\mu_{3u}\frac{(n-1)(n-2)}{n^{2}}+\mu_{3v}\Big(\frac{1}{J_{i}^{2}}-\frac{3}{nJ_{i}^{2}}+\frac{3}{n^{2}J_{i}^{2}}-\frac{1}{n^{3}}\sum_{i^{\prime}=1}^{n}\frac{1}{J_{i^{\prime}}^{2}}\Big)$
$\quad+\mu_{3w}\Bigg(\Big(1-\frac{1}{n}\Big)^{3}\frac{1}{J_{i}^{3}}\sum_{j=1}^{J_{i}}\frac{1}{K_{ij}^{2}}+\Big(-\frac{1}{n}\Big)^{3}\sum_{i^{\prime}\neq i}\frac{1}{J_{i^{\prime}}^{3}}\sum_{j=1}^{J_{i^{\prime}}}\frac{1}{K_{i^{\prime}j}^{2}}\Bigg).$
Similarly, using between-$i$ variation and observation-averages
$\bar{y}_{i}^{\mathrm{o}}-\bar{y}^{\mathrm{o}}=[u_{i}-\bar{u}^{\mathrm{o}}]+[\bar{v}_{i}^{\mathrm{o}}-\bar{v}^{\mathrm{o}}]+[\bar{w}_{i}^{\mathrm{o}}-\bar{w}^{\mathrm{o}}]$ $=\Big[\Big(1-\frac{K_{i}}{N}\Big)u_{i}-\frac{1}{N}\sum_{i^{\prime}\neq i}K_{i^{\prime}}u_{i^{\prime}}\Big]+\Big[\Big(1-\frac{K_{i}}{N}\Big)\frac{1}{K_{i}}\sum_{j=1}^{J_{i}}K_{ij}v_{ij}-\frac{1}{N}\sum_{i^{\prime}\neq i}\sum_{j=1}^{J_{i^{\prime}}}K_{i^{\prime}j}v_{i^{\prime}j}\Big]$
$\quad+\Big[\Big(1-\frac{K_{i}}{N}\Big)\frac{1}{K_{i}}\sum_{j=1}^{J_{i}}\sum_{k=1}^{K_{ij}}w_{ijk}-\frac{1}{N}\sum_{i^{\prime}\neq i}\sum_{j=1}^{J_{i^{\prime}}}\sum_{k=1}^{K_{i^{\prime}j}}w_{i^{\prime}jk}\Big].$
Applying Lemma A.1 and using independence gives
$\mathbb{E}[(\bar{y}_{i}^{\mathrm{o}}-\bar{y}^{\mathrm{o}})^{2}]$ $=\mu_{2u}\Big(1-\frac{2K_{i}}{N}+\frac{1}{N^{2}}\sum_{i^{\prime}=1}^{n}K_{i^{\prime}}^{2}\Big)+\mu_{2v}\Bigg(\Big(1-\frac{K_{i}}{N}\Big)^{2}\frac{1}{K_{i}^{2}}\sum_{j=1}^{J_{i}}K_{ij}^{2}+\frac{1}{N^{2}}\sum_{i^{\prime}\neq i}\sum_{j=1}^{J_{i^{\prime}}}K_{i^{\prime}j}^{2}\Bigg)+\mu_{2w}\Big(\frac{1}{K_{i}}-\frac{1}{N}\Big),$
$\mathbb{E}[(\bar{y}_{i}^{\mathrm{o}}-\bar{y}^{\mathrm{o}})^{3}]$ $=\mu_{3u}\Big(1-\frac{3K_{i}}{N}+\frac{3K_{i}^{2}}{N^{2}}-\frac{1}{N^{3}}\sum_{i^{\prime}=1}^{n}K_{i^{\prime}}^{3}\Big)+\mu_{3v}\Bigg(\Big(1-\frac{K_{i}}{N}\Big)^{3}\frac{1}{K_{i}^{3}}\sum_{j=1}^{J_{i}}K_{ij}^{3}+\Big(-\frac{1}{N}\Big)^{3}\sum_{i^{\prime}\neq i}\sum_{j=1}^{J_{i^{\prime}}}K_{i^{\prime}j}^{3}\Bigg)$
$\quad+\mu_{3w}\frac{(N-K_{i})(N-2K_{i})}{N^{2}K_{i}^{2}}.$